# A highly efficient single-loop smoothing damped Newton method for large-scale bilevel hyperparameter optimization of SVC

### v1

@@ 2 Bilevel hyperparameter optimization for SVC
$A:={\rm Diag}\left(A^{1},\ A^{2},\cdots,A^{n}\right)\in\mathbb{R}^{Tm_{1}\times Tn},\ B:={\rm Diag}\left(B^{1},\ B^{2},\cdots,B^{n}\right)\in\mathbb{R}^{Tm_{1}\times Tn},$
$A^{t}:=\left[\begin{array}[]{c}y_{t_{1}}x_{t_{1}}^{\top}\\ \vdots\\ y_{t_{m_{1}}}x_{t_{m_{1}}}^{\top}\end{array}\right]\in\mathbb{R}^{m_{1}\times n},\ B^{t}:=\begin{bmatrix}y_{t_{m_{1}+1}}x_{t_{m_{1}+1}}^{\top}\\ \vdots\\ y_{t_{l_{1}}}x_{t_{l_{1}}}^{\top}\end{bmatrix}\in\mathbb{R}^{m_{2}\times n},\ t\in[T],\$
$\small\zeta:=\left[\begin{array}[]{l}\zeta^{1}\\ \zeta^{2}\\ \vdots\\ \zeta^{T}\end{array}\right]\in\mathbb{R}^{Tm_{1}},\ z:=\left[\begin{array}[]{l}z^{1}\\ z^{2}\\ \vdots\\ z^{T}\end{array}\right]\in\mathbb{R}^{Tm_{1}},\ \alpha:=\left[\begin{array}[]{l}\alpha^{1}\\ \alpha^{2}\\ \vdots\\ \alpha^{T}\end{array}\right]\in\mathbb{R}^{Tm_{2}},\ \xi:=\left[\begin{array}[]{l}\xi^{1}\\ \xi^{2}\\ \vdots\\ \xi^{T}\end{array}\right]\in\mathbb{R}^{Tm_{2}}.$
(2) can be written into the following compact form:
$\begin{array}[]{cl}\min\limits_{v\in\mathbb{R}^{m+1}}&f(v)\\ \hbox{s.t.}&0\leq H(v)\perp G(v)\geq 0,\end{array}$ (MPEC)
where $m=2T(m_{1}+m_{2}),\ f(v):=\frac{1}{Tm_{1}}\mathbf{1}^{\top}\zeta$, and $G(v)\in\mathbb{R}^{m},\ H(v)\in\mathbb{R}^{m},\ v\in\mathbb{R}^{m+1}$ are defined by
@@
$G(v):=\left[\begin{array}[]{c}\zeta\\ z\\ \alpha\\ \xi\end{array}\right]:=L^{G}v,\ H(v):=\left[\begin{array}[]{c}AB^{\top}\alpha+z\\ \mathbf{1}-\zeta\\ BB^{\top}\alpha-\mathbf{1}+\xi\\ C\mathbf{1}-\alpha\end{array}\right]:=L^{H}v+b^{H},\ v:=\left[\begin{array}[]{c}C\\ \zeta\\ z\\ \alpha\\ \xi\end{array}\right].$ (3)
Here, $L^{G},L^{H}\in\mathbb{R}^{m\times(m+1)}$, and $b^{H}\in\mathbb{R}^{m}$ are given by
$L^{G}=\left[\begin{array}[]{ccccc}\mathbf{0}&\mathcal{I}&\mathbf{0}&\mathbf{0}&\mathbf{0}\\ \mathbf{0}&\mathbf{0}&\mathcal{I}&\mathbf{0}&\mathbf{0}\\ \mathbf{0}&\mathbf{0}&\mathbf{0}&\mathcal{I}&\mathbf{0}\\ \mathbf{0}&\mathbf{0}&\mathbf{0}&\mathbf{0}&\mathcal{I}\\ \end{array}\right],\ L^{H}=\left[\begin{array}[]{ccccc}\mathbf{0}&\mathbf{0}&\mathcal{I}&AB^{\top}&\mathbf{0}\\ \mathbf{0}&-\mathcal{I}&\mathbf{0}&\mathbf{0}&\mathbf{0}\\ \mathbf{0}&\mathbf{0}&\mathbf{0}&BB^{\top}&\mathcal{I}\\ \mathbf{1}&\mathbf{0}&\mathbf{0}&-\mathcal{I}&\mathbf{0}\\ \end{array}\right],\ b^{H}=\left[\begin{array}[]{c}\mathbf{0}\\ \mathbf{1}\\ \mathbf{-1}\\ \mathbf{0}\\ \end{array}\right].$ (4)
Specifically, one can notice that the number of components in $v\in\mathbb{R}^{n+1}$ and the number of complementarity constraints are $m+1$ and $m$, respectively, where $m=2T(m_{1}+m_{2})$, $m_{1}$ and $m_{2}$ are the number of training data and validation data, respectively. Therefore, when the number of cross validation or the number of dataset increases, it will lead to the increase of the scale of (MPEC). Solving such a large-scale bilevel hyperparameter selection problem for SVC is extremely challenging, in which the cross-validation process increases the scale of the problem exponentially.
Keeping in mind that our aim is to design an efficient numerical algorithm to solve (MPEC), which is single-loop, and also software-independent, the complementarity constraints are the first issue that we want to tackle. In the coming section, we will smooth them and investigate related properties.
3 Smoothing optimization problem of (MPEC)
In this section, we will consider the smoothing problem of (MPEC) by introducing smoothing functions to deal with the complementarity constraints in (MPEC). We will also investigate the relation of the smoothing problem and (MPEC).
In (MPEC), $0\leq H(v)\perp G(v)\geq 0$ are linear complementarity constraints. Therefore, from the optimization point of view, there are various methods to tackle complementarity constraints to their smoothing equivalent system to make the computation process more efficient [6, 21, 23, 24]. Different smoothing functions have been proposed and studied [3, 4, 14]. Here we choose the popular Fischer-Burmeister smoothing function [3], which is defined by
Then $\phi_{\epsilon}(a,b)=0$ if and only if $a>0,b>0,ab=\frac{\epsilon^{2}}{2}$. For $s\in\mathbb{R}^{m},\ t\in\mathbb{R}^{m}$, define
$\Phi_{\epsilon}(s,t)=\left[\begin{array}[]{c}\phi_{\epsilon}(s_{1},t_{1})\\ \vdots\\ \phi_{\epsilon}(s_{m},t_{m})\end{array}\right]\in\mathbb{R}^{m}.$
Applying $\Phi(\cdot,\cdot)$ to the complementarity constraints in (MPEC), we reach the smoothing version of (MPEC) as follows ($\epsilon>0$), which is a essentially a nonlinear programming problem with equilibrium constraints:
$\begin{array}[]{cl}\min\limits_{v\in\mathbb{R}^{m+1}}&f(v)\\ \hbox{s.t.}&\Phi_{\epsilon}(G(v),H(v))=0.\end{array}$ (NLPEC)
Here we would like to make a comparison between (MPEC) and (NLPEC). In terms of variables, the two problems have the same scale of variables, which is $m+1$. However, the number of constraints in (NLPEC) is only one third of that in (MPEC), which is much smaller. Moreover, the constraints in (NLPEC) are equality constraints, which are much easier to deal with in designing algorithms.
@@ 3 Smoothing optimization problem of (MPEC)
Before focusing on solving the smoothing problem (NLPEC), we need to investigate the feasible region of (NLPEC) and (MPEC). Due to the smoothing function in (NLPEC) and the positive smoothing parameter $\epsilon$, the feasible region of (NLPEC) is no longer the same as that of (MPEC). Therefore, what is the relation of feasible sets of (NLPEC) and (MPEC), as $\epsilon\rightarrow 0^{+}$? We close this section by addressing this question below.
Denote the feasible sets of (MPEC) and (NLPEC) by
@@
To answer the above question, we need to reformulate (MPEC) and (NLPEC) as the following equivalent form
$\begin{array}[]{cl}\min\limits_{v\in\mathbb{R}^{m+1},p\in\mathbb{R}^{m},q\in\mathbb{R}^{m}}&f(v)\\ \hbox{s.t.}&G(v)=p,\ H(v)=q,\\ &0\leq G(v)\perp H(v)\geq 0,\end{array}$ (5)
and
$\begin{array}[]{cl}\min\limits_{v\in\mathbb{R}^{m+1},p\in\mathbb{R}^{m},q\in\mathbb{R}^{m}}&f(v)\\ \hbox{s.t.}&G(v)=p,\ H(v)=q,\\ &\Phi_{\epsilon}(G(v),H(v))=0.\end{array}$ (6)
@@
$\Gamma^{0}:=\Gamma^{=}\cap\Gamma^{\perp},\ \Gamma(\epsilon):=\Gamma^{=}\cap\Gamma^{\perp}(\epsilon),$ (7)
$\Gamma^{=}:=\{(v,p,q)\mid G(v)=p,\ H(v)=q\},$
$\Gamma^{\perp}:=\{(v,p,q)\mid 0\leq G(v)\perp H(v)\geq 0\},$
$\Gamma^{\perp}(\epsilon):=\{(v,p,q)\mid\Phi_{\epsilon}(G(v),H(v)=0\}.$
To study the relationship between $\Gamma(\epsilon)$ and $\Gamma^{0}$, we first consider the relationship between $\underset{\epsilon\searrow 0}{\lim}\sup\Gamma(\epsilon)$ and $\Gamma^{0}$ in the following result.
@@
Let $\Gamma(\epsilon)$ and $\Gamma^{0}$ be defined in (7), it holds that
$\underset{\epsilon\searrow 0}{\lim}\sup\Gamma(\epsilon)\subseteq\Gamma^{0}.$
@@
For any $(v,p,q)\in\underset{\epsilon\searrow 0}{\lim}\sup\Gamma(\epsilon)$, by definition, there exists $\epsilon^{k}\searrow 0$ and $(v^{k},p^{k},q^{k})\to(v,p,q)$ with $(v^{k},p^{k},q^{k})\in\Gamma(\epsilon^{k})$. Next, we need to show that $(v,p,q)\in\Gamma^{0}$. To that end, by taking limits, we obtain that
@@
Moreover, we have
$\underset{k\to\epsilon}{\lim}\Phi_{\epsilon}(p^{k},q^{k})=\underset{k\to\epsilon}{\lim}\frac{(\epsilon^{k})^{2}}{2}=0,$
implying that $0\leq G(v)\perp H(v)\geq 0$. This gives that $(v,p,q)\in\Gamma^{0}$. The proof is finished. ∎
Next, to show that $\Gamma^{0}\subseteq\underset{\epsilon\searrow 0}{\lim}\inf\Gamma(\epsilon)$, we define the following sets for each $(v,p,q)\in\Gamma^{0}$:
$I_{0+}=\{i\in m\mid p_{i}=0,\ q_{i}>0\},$
$I_{+0}=\{i\in m\mid p_{i}>0,\ q_{i}=0\},$
$I_{00}=\{i\in m\mid p_{i}=0,\ q_{i}=0\},$
and define $\theta_{\epsilon}(t)=\left(\theta_{\epsilon,1}(t_{1}),\theta_{\epsilon,2}(t_{2}),\cdots,\theta_{\epsilon,m}(t_{m})\right)^{\top}$ by
$\theta_{\epsilon,j}(t_{j}):=\begin{cases}\frac{\epsilon^{2}}{2t_{j}}-q_{j},&\ j\in I_{0+},\\ \frac{\epsilon^{2}}{2\left(p_{j}+t_{j}\right)},&\ j\in I_{+0},\\ \ \ \ \frac{\epsilon^{2}}{2t_{j}},&\ j\in I_{00}.\end{cases}$
@@
$\begin{cases}\Delta p^{k}\geq 0,\theta_{\epsilon^{k}}(\Delta p^{k})\geq 0,(\Delta C^{k},\Delta p^{k},\theta_{\epsilon^{k}}(\Delta p^{k}))\to 0,\\ L^{H}_{(:,1)}\Delta C^{k}+L^{H}_{(:,2:n)}\Delta p^{k}=\theta_{\epsilon^{k}}(\Delta p^{k}).\end{cases}$
@@
Under Assumption 3.2, it holds that
$\Gamma^{0}\subseteq\underset{\epsilon\searrow 0}{\lim}\inf\Gamma(\epsilon).$
@@
To proof the result, we need to show that for any $(v,p,q)\in\Gamma^{0}$, $(v,p,q)\in\underset{\epsilon\searrow 0}{\lim}\inf\Gamma(\epsilon)$. That is, for any $\epsilon^{k}\searrow 0$, there exists a sequence $(v^{k},p^{k},q^{k})$, such that
$\lim\limits_{k\to\infty}(v^{k},p^{k},q^{k})\to(v,p,q),\ {\rm and}\ (v^{k},p^{k},q^{k})\in\Gamma(\epsilon^{k}),$
which is,
$\lim\limits_{k\to\infty}(v^{k},p^{k},q^{k})=(v,p,q),\ H(v^{k})=q^{k},\ G(v^{k})=p^{k},\ \Phi_{\epsilon}(p^{k},q^{k})=\frac{(\epsilon^{k})^{2}}{2}.$ (8)
@@
$(\Delta v^{k},\Delta p^{k},\Delta q^{k})\to(0,0,0),\ \epsilon^{k}\to 0,$ (9d)
@@
$\Delta p^{k}\geq 0,\ \Delta q^{k}\geq 0,\ (p_{j}+\Delta p^{k}_{j})(q_{j}+\Delta q^{k}_{j})=\frac{(\epsilon^{k})^{2}}{2}.$
Recall the definition of $G(\cdot)$ and $H(\cdot)$ in (3), the above system reduces to the following
$\begin{cases}L^{H}(v+\Delta v^{k})+b^{H}=q+\Delta q^{k},\\ L^{G}(v+\Delta v^{k})=p+\Delta p^{k},\\ \Delta p^{k}\geq 0,\ \Delta q^{k}\geq 0,\ (p_{j}+\Delta p^{k}_{j})(q_{j}+\Delta q^{k}_{j})=\frac{(\epsilon^{k})^{2}}{2},\\ (\Delta v^{k},\Delta p^{k},\Delta q^{k})\to(0,0,0),\ \epsilon^{k}\to 0.\end{cases}$ (10)
@@
$\begin{cases}L^{H}v+b^{H}=q,\\ L^{G}v=p,\\ p_{I_{0+}}=0,\ p_{I_{+0}}>0,\ p_{I_{00}}=0,\\ q_{I_{0+}}>0,\ q_{I_{+0}}=0,\ q_{I_{00}}=0.\end{cases}$
@@
$=$ $\begin{cases}q_{j}\Delta p_{j}^{k}+\Delta p_{j}^{k}\Delta q_{j}^{k},&i\in I_{0+},\\ p_{j}\Delta q_{j}^{k}+\Delta p_{j}^{k}\Delta q_{j}^{k},&i\in I_{+0},\\ \Delta p_{j}^{k}\Delta q_{j}^{k},&i\in I_{00}.\end{cases}$

### v3

@@ 3 Solution Method
$\begin{array}[]{cc}\hat{\gamma_{i}}=\begin{cases}\hat{\lambda}_{i}>0,&i\in\hat{I}_{+}\cap(I_{0+}\cup I_{00}^{2G}),\\ \hat{\lambda}_{i}a_{i}>0,&i\in\hat{I}_{+}\cap I_{00}^{3G},\\ \hat{\lambda}_{i}<0,&i\in\hat{I}_{-}\cap(I_{0+}\cup I_{00}^{2G}),\\ \hat{\lambda}_{i}a_{i}<0,&i\in\hat{I}_{-}\cap I_{00}^{3G},\\ 0,&i\in I_{+0}\cup I_{00}^{1G}\cup\hat{I}_{0},\end{cases}&\hat{\nu_{i}}=\begin{cases}\hat{\lambda}_{i}>0,&i\in\hat{I}_{+}\cap(I_{+0}\cup I_{00}^{2H}),\\ \hat{\lambda}_{i}b_{i}>0,&i\in\hat{I}_{+}\cap I_{00}^{3H},\\ \hat{\lambda}_{i}<0,&i\in\hat{I}_{-}\cap(I_{+0}\cup I_{00}^{2H}),\\ \hat{\lambda}_{i}b_{i}<0,&i\in\hat{I}_{-}\cap I_{00}^{3H},\\ 0,&i\in I_{0+}\cup I_{00}^{1H}\cup\hat{I}_{0}.\end{cases}\end{array}$
Therefore, taking limits for $t\to+\infty$ in (4) gives
$0=-\sum_{i\in I_{0+}\cup I_{00}^{2G}\cup I_{00}^{3G}}\hat{\gamma_{i}}\nabla G_{i}(v^{*})-\sum_{i\in I_{+0}\cup I_{00}^{2H}\cup I_{00}^{3H}}\hat{\nu_{i}}\nabla H_{i}(v^{*}),$
where $\hat{\gamma_{i}}\neq 0,i\in i\in I_{0+}\cup I_{00}^{2G}\cup I_{00}^{3G};\hat{\nu_{i}}\neq 0,i\in I_{+0}\cup I_{00}^{2H}\cup I_{00}^{3H}$. Therefore, it is a contradiction to the prerequisite that MPEC-LICQ holds at $v^{*}$, since $I_{00}^{2G}\cup I_{00}^{3G}\subseteq I_{00}$ and $I_{00}^{2H}\cup I_{00}^{3H}\subseteq I_{00}$. Thus, we may assume without loss of generality that one of the accumulation points of $\left\{\lambda^{k}\right\}$ is $\lambda^{*}$. The continuous differentiability of $f,G,H$ implies
$\textbf{0}=\nabla f(v^{*})-\sum_{i=1}^{m}\gamma_{i}^{*}\nabla G_{i}(v^{*})-\sum_{i=1}^{m}\nu_{i}^{*}\nabla H_{i}(v^{*}),$ (5)
where $\gamma_{i}^{*}=\lambda^{*}_{i}a_{i},$, $\nu_{i}^{*}=\lambda^{*}_{i}b_{i}$.
By defining $_{+}^{*}, II_{-}^{*}$ and $I_{0}^{*}$ on $\lambda^{*}$ similar to that of $\hat{I}_{+}$, $\hat{I}_{-}$ and $\hat{I}_{0}$, one can obtain the following
$\begin{array}[]{c}\gamma_{i}^{*}=\begin{cases}\lambda_{i}^{*}>0,&i\in I_{+}^{*}\cap(I_{0+}\cup I_{00}^{2G}),\\ \lambda_{i}^{*}a_{i}>0,&i\in I_{+}^{*}\cap I_{00}^{3G},\\ \lambda_{i}^{*}<0,&i\in I_{-}^{*}\cap(I_{0+}\cup I_{00}^{2G}),\\ \lambda_{i}^{*}a_{i}<0,&i\in I_{-}^{*}\cap I_{00}^{3G},\\ 0,&i\in I_{+0}\cup I_{00}^{1G}\cup I_{0}^{*},\end{cases}\ \nu_{i}^{*}=\begin{cases}\lambda_{i}^{*}>0,&i\in I_{+}^{*}\cap(I_{+0}\cup I_{00}^{2H}),\\ \lambda_{i}^{*}b_{i}>0,&i\in I_{+}^{*}\cap I_{00}^{3H},\\ \lambda_{i}^{*}<0,&i\in I_{-}^{*}\cap(I_{+0}\cup I_{00}^{2H}),\\ \lambda_{i}^{*}b_{i}<0,&i\in I_{-}^{*}\cap I_{00}^{3H},\\ 0,&i\in I_{0+}\cup I_{00}^{1H}\cup I_{0}^{*}.\end{cases}\end{array}$
It holds that
$0\leq\gamma_{i}^{*}\nu_{i}^{*}=\begin{cases}(\lambda_{i}^{*})^{2},&i\in(I_{+}^{*}\cup I_{-}^{*})\cap(I_{0+}\cup I_{00}^{2G})\cap(I_{+0}\cup I_{00}^{2H}),\\ (\lambda_{i}^{*})^{2}a_{i},&i\in(I_{+}^{*}\cup I_{-}^{*})\cap I_{00}^{3G}\cap(I_{+0}\cup I_{00}^{2H}),\\ (\lambda_{i}^{*})^{2}b_{i},&i\in(I_{+}^{*}\cup I_{-}^{*})\cap I_{00}^{3H}\cap(I_{0+}\cup I_{00}^{2G}),\\ (\lambda_{i}^{*})^{2}a_{i}b_{i},&i\in(I_{+}^{*}\cup I_{-}^{*})\cap I_{00}^{3G}\cap I_{00}^{3H},\\ 0,&{\rm otherwise.}\end{cases}$
Overall, based on the definition of C-stationary point, $v^{*}$ is a C-stationary point of (MPEC). The proof is finished. ∎
Based on Theorem 3.1, to guarantee the convergence to a C-stationary point,we need the MPEC-LICQ at accumulation point $v^{*}$, as well as the LICQ at $v^{t}$ for each $t$. The MPEC-LICQ is guaranteed when $\left|I_{GH}\right|=1$ by the following theorem proposed in [22]. The fulfillment of LICQ at each $v^{t}$ will be addressed in Section 5.
Theorem 3.2.
Let $v^{*}$ be a feasible point of problem (MPEC).
  • $\left(i\right)$
    If $\left|I_{GH}\right|>1$, then MPEC-LICQ fails at $v^{*}$.
  • $\left(ii\right)$
    If $\left|I_{GH}\right|=0$, then MPEC-LICQ holds at $v^{*}$.
  • $\left(iii\right)$
    If $\left|I_{GH}\right|=1$, $(BB^{\top})_{(\Lambda_{3}^{+},\Lambda_{3}^{+})}$ is positive definite at $v^{*}$ and $\Gamma_{sub}\neq 0$, then MPEC-LICQ holds at $v^{*}$, where
    $\footnotesize\Gamma_{\mathrm{sub}}=\begin{bmatrix}(BB^{\top})_{(I_{GH_{4}},\Lambda_{3}^{c}\cup\Lambda_{u})}\,\mathbf{1}-(BB^{\top})_{(I_{GH_{4}},\Lambda_{3}^{+})}\,(BB^{\top})_{(\Lambda_{3}^{+},\Lambda_{3}^{+})}^{-1}\,(BB^{\top})_{(\Lambda_{3}^{+},\Lambda_{3}^{c}\cup\Lambda_{u})}\,\mathbf{1}\\ (BB^{\top})_{(I_{GH_{3}},\Lambda_{3}^{c}\cup\Lambda_{u})}\,\mathbf{1}-(BB^{\top})_{(I_{GH_{3}},\Lambda_{3}^{+})}\,(BB^{\top})_{(\Lambda_{3}^{+},\Lambda_{3}^{+})}^{-1}\,(BB^{\top})_{(\Lambda_{3}^{+},\Lambda_{3}^{c}\cup\Lambda_{u})}\,\mathbf{1}\end{bmatrix}$
    and
@@
    $\begin{cases}\Lambda_{3}^{+}:=\{i\in Q_{l}\ |\ 0<\alpha_{i}<C,\ (BB^{\top}\alpha-\mathbf{1}+\xi)_{i}=0,\ \xi_{i}=0\},\\ \Lambda_{3}^{c}:=\left\{i\in Q_{l}\ |\ \alpha_{i}=C,(BB^{\top}\alpha-\mathbf{1}+\xi)_{i}=0,\ \xi_{i}=0\right\},\\ \Lambda_{u}:=\left\{i\in Q_{l}\ |\ \alpha_{i}=C,\ (BB^{\top}\alpha-\mathbf{1}+\xi)_{i}=0,\ \xi_{i}>0\right\},\\ I_{GH_{3}}:=\{i\in Q_{l}\ |\ \alpha_{i}=0,\ (BB^{\top}\alpha-\mathbf{1}+\xi)_{i}=0\},\\ I_{GH_{4}}:=\{i\in Q_{l}\ |\ \xi_{i}=0,\ C-\alpha_{i}=0\},\\ Q_{l}:=\{1,2,\cdots,Tm_{2}\}.\end{cases}$
4 Feasible Sets of (NLPϵ) and (MPEC)
@@
Before focusing on solving the smoothing subproblem (NLPϵ), we need to investigate the feasible region of (NLPϵ) and (MPEC). Due to the smoothing function in (NLPϵ) and the positive smoothing parameter $\epsilon$, the feasible region of (NLPϵ) is no longer the same as that of (MPEC). Therefore, what is the relation of feasible sets of (NLPϵ) and (MPEC), as $\epsilon\searrow 0$? We address this question below.
Denote the feasible sets of (MPEC) and (NLPϵ) by
@@
Consider the following equivalent form of (MPEC) and (NLPϵ)
$\begin{array}[]{cl}\min\limits_{v\in\mathbb{R}^{m+1},p\in\mathbb{R}^{m},q\in\mathbb{R}^{m}}&f(v)\\ \hbox{s.t.}&G(v)=p,\ H(v)=q,\ 0\leq p\perp q\geq 0,\end{array}$ (8)
$\begin{array}[]{cl}\min\limits_{v\in\mathbb{R}^{m+1},p\in\mathbb{R}^{m},q\in\mathbb{R}^{m}}&f(v)\\ \hbox{s.t.}&G(v)=p,\ H(v)=q,\ \Phi_{\epsilon}(p,q)=0.\end{array}$ (11)
@@
$\Gamma^{0}=\Gamma^{=}\cap\Gamma^{\perp},\ \Gamma(\epsilon)=\Gamma^{=}\cap\Gamma^{\perp}(\epsilon),\ \Gamma^{=}=\{(v,p,q)\mid G(v)=p,\ H(v)=q\},$ (12)
$\Gamma^{\perp}=\{(v,p,q)\mid 0\leq p\perp q\geq 0\},\ \Gamma^{\perp}(\epsilon)=\{(v,p,q)\mid\Phi_{\epsilon}(p,q)=0\}.$
We first address the relationship between $\underset{\epsilon\searrow 0}{\lim}\sup\Gamma(\epsilon)$ and $\Gamma^{0}$ in the following lemma.
@@
Let $\Gamma(\epsilon)$ and $\Gamma^{0}$ be defined in (12), it holds that $\underset{\epsilon\searrow 0}{\lim}\sup\Gamma(\epsilon)\subseteq\Gamma^{0}.$
@@
For any $(v,p,q)\in\underset{\epsilon\searrow 0}{\lim}\sup\Gamma(\epsilon)$, by definition, there exists $\epsilon_{k}\searrow 0$ and $(v^{k},p^{k},q^{k})\to(v,p,q)$ with $(v^{k},p^{k},q^{k})\in\Gamma(\epsilon_{k})$. Next, we need to show that $(v,p,q)\in\Gamma^{0}$. To that end, by taking limits, we obtain that
@@
Moreover, we have $\underset{k\to\epsilon}{\lim}\Phi_{\epsilon}(p^{k},q^{k})=\underset{k\to\epsilon}{\lim}\frac{(\epsilon_{k})^{2}}{2}=0,$ implying that $0\leq G(v)\perp H(v)\geq 0$. This gives that $(v,p,q)\in\Gamma^{0}$. The proof is finished. ∎
To show that $\Gamma^{0}\subseteq\underset{\epsilon\searrow 0}{\lim}\inf\Gamma(\epsilon)$, we define the following sets for each $(v,p,q)\in\Gamma^{0}$: $I_{0+}^{\Gamma}=\{i\in m\mid p_{i}=0,\ q_{i}>0\}$, $I_{+0}^{\Gamma}=\{i\in m\mid p_{i}>0,\ q_{i}=0\}$, $I_{00}^{\Gamma}=\{i\in m\mid p_{i}=0,\ q_{i}=0\}$, and define $\theta_{\epsilon}(\iota)=\left(\theta_{\epsilon,1}(\iota_{1}),\theta_{\epsilon,2}(\iota_{2}),\cdots,\theta_{\epsilon,m}(\iota_{m})\right)^{\top}$ by
$\theta_{\epsilon,j}(\iota_{j}):=\begin{cases}\frac{\epsilon^{2}}{2\iota_{j}}-q_{j},&\ j\in I_{0+}^{\Gamma},\\ \frac{\epsilon^{2}}{2\left(p_{j}+\iota_{j}\right)},&\ j\in I_{+0}^{\Gamma},\\ \ \ \ \frac{\epsilon^{2}}{2\iota_{j}},&\ j\in I_{00}^{\Gamma}.\end{cases}$
@@
$\begin{cases}\Delta p^{k}\geq 0,\theta_{\epsilon_{k}}(\Delta p^{k})\geq 0,(\Delta C^{k},\Delta p^{k},\theta_{\epsilon_{k}}(\Delta p^{k}))\to 0,\\ L^{H}_{(:,1)}\Delta C^{k}+L^{H}_{(:,2:n)}\Delta p^{k}=\theta_{\epsilon_{k}}(\Delta p^{k}).\end{cases}$
@@
Under Assumption 4.2, it holds that $\Gamma^{0}\subseteq\underset{\epsilon\searrow 0}{\lim}\inf\Gamma(\epsilon).$
@@
To proof the result, we need to show that for any $(v,p,q)\in\Gamma^{0}$, $(v,p,q)\in\underset{\epsilon\searrow 0}{\lim}\inf\Gamma(\epsilon)$. That is, for any $\epsilon_{k}\searrow 0$, there exists a sequence $(v^{k},p^{k},q^{k})$, such that $\lim\limits_{k\to\infty}(v^{k},p^{k},q^{k})\to(v,p,q),\ {\rm and}\ (v^{k},p^{k},q^{k})\in\Gamma(\epsilon_{k})$, which is,
$\lim\limits_{k\to\infty}(v^{k},p^{k},q^{k})=(v,p,q),\ H(v^{k})=q^{k},\ G(v^{k})=p^{k},\ \Phi_{\epsilon}(p^{k},q^{k})=\frac{(\epsilon_{k})^{2}}{2}.$ (13)
@@
$(\Delta v^{k},\Delta p^{k},\Delta q^{k})\to(0,0,0),\ \epsilon_{k}\to 0,$ (14d)
@@
$\Delta p^{k}\geq 0,\ \Delta q^{k}\geq 0,\ (p_{j}+\Delta p^{k}_{j})(q_{j}+\Delta q^{k}_{j})=\frac{(\epsilon_{k})^{2}}{2}.$
Recall the definition of $G(\cdot)$ and $H(\cdot)$ in (3). $G(v)$ and $H(v)$ can be written as $G(v)=L^{G}v$, and $H(v)=L^{H}v+b^{H}$. The above system reduces to the following
$\begin{cases}L^{H}(v+\Delta v^{k})+b^{H}=q+\Delta q^{k},\\ L^{G}(v+\Delta v^{k})=p+\Delta p^{k},\\ \Delta p^{k}\geq 0,\ \Delta q^{k}\geq 0,\ (p_{j}+\Delta p^{k}_{j})(q_{j}+\Delta q^{k}_{j})=\frac{(\epsilon_{k})^{2}}{2},\\ (\Delta v^{k},\Delta p^{k},\Delta q^{k})\to(0,0,0),\ \epsilon_{k}\to 0.\end{cases}$ (15)
@@
$\begin{cases}L^{H}v+b^{H}=q,\\ L^{G}v=p,\\ p_{I_{0+}^{\Gamma}}=0,\ p_{I_{+0}^{\Gamma}}>0,\ p_{I_{00}^{\Gamma}}=0,\\ q_{I_{0+}^{\Gamma}}>0,\ q_{I_{+0}^{\Gamma}}=0,\ q_{I_{00}^{\Gamma}}=0.\end{cases}$
@@
$=$ $\begin{cases}q_{j}\Delta p_{j}^{k}+\Delta p_{j}^{k}\Delta q_{j}^{k},&j\in I_{0+}^{\Gamma},\\ p_{j}\Delta q_{j}^{k}+\Delta p_{j}^{k}\Delta q_{j}^{k},&j\in I_{+0}^{\Gamma},\\ \Delta p_{j}^{k}\Delta q_{j}^{k},&j\in I_{00}^{\Gamma}.\end{cases}$